Question 5 (5 points) Expedia would like to assess the hypothesis that the proportion of United Airline flights that arrive on-time is less than 0.80. A random sample of 110 United Airline flights found that 82 arrived on-time. Expedia would like to set α = 0.02. The p-value for this hypothesis test would be ________.

Question 5 options:

a) 0.0322

b) 0.0448

c) 0.0764

d) 0.0912
Answer #1

Solution :

This is the left tailed test .

The null and alternative hypothesis is

H0 : p = 0.80

Ha : p < 0.80

\hat p  = x / n = 82 / 110 = 0.7455

P0 = 0.80

1 - P0 = 1 - 0.80 = 0.20

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.7455 -0.80/ [\sqrt(0.80 * 0.20) / 110]

Test statistic = z = -1.43

P(z < -1.43)

P-value = 0.0764

correct option is = c